OpenStudy (anonymous):

You are starting a dog-walking business to earn some extra money around your neighborhood. The function r = 10d – d2 gives you the estimated revenue r, that you can expect to earn each day from walking d number of dogs. From past experience, you know that the number of dogs d that you will have the opportunity to walk varies with the price p that you charge per dog given by the equation d = 10 – p.

OpenStudy (anonymous):

a. Describe the graph of each function and use the information above to find the greatest amount of revenue that you can expect to receive from walking dogs. It may help you to graph each function separately and label each axis. b. In order to get the maximum amount of revenue that you just found, what price should you charge to walk each dog?

OpenStudy (anonymous):



OpenStudy (anonymous):

Revenue R(d) = – d² + 10d Max value happens at d = 10/2 = 5 dogs -> R (5) = ....?

OpenStudy (anonymous):

@Student477777 Can you finish it and post your answer so I can double check?

OpenStudy (anonymous):

Im still trying to understand Ill post what I come up with

OpenStudy (anonymous):

Yeah nope I dont get it xD